# Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models

### v1

@@ B.1 Category and Inflection Distribution
Table 3: Distribution of word categories in the dataset
Table 4 presents the distribution of inflection categories across the dataset. Singular nouns are the most frequent form (36.2%), while adjectival forms like comparative and superlative are relatively rare (0.7% and 0.5% respectively).
@@
Table 4: Distribution of inflection categories in the dataset
B.2 Sentence Statistics
Sentences in the corpus are relatively concise, with an average length of 6.5 words and a median of 5 words. The minimum number of words is 1 and the maximum is 40. This relatively short average length is ideal for our analysis, as it reduces the complexity of contextual dependencies while still providing sufficient context for meaningful representation.
Figure 6 shows the distribution of sentence lengths.
Figure 6: Distribution of sentence lengths in the dataset
B.3 Tokenization Statistics
@@
An important consideration for our analysis is how different models tokenize the words in our dataset. Table 6 shows tokenization statistics across the models we analyze. Encoder-only models like BERT and DeBERTa tend to split words into more tokens than decoder-only models like GPT-2 and Qwen2, which may affect how information is encoded across layers.
Avg. tokens Med. tokens Max tokens Percent multitoken
Model per word per word per word
BERT-base 3.11 3.0 8.0 100.0
BERT-large 3.11 3.0 8.0 100.0
DeBERTa-v3-large 3.03 3.0 6.0 100.0
GPT-2-small 1.52 1.0 5.0 42.25
GPT-2-large 1.52 1.0 5.0 42.25
GPT-2-xl 1.52 1.0 5.0 42.25
Pythia-6.9B 1.48 1.0 5.0 39.1
Pythia-6.9B-Tulu 1.48 1.0 5.0 39.1
OLMo2-7B 1.43 1.0 4.0 35.9
OLMo2-7B-Instruct 1.43 1.0 4.0 35.9
Gemma2-2B 2.19 2.0 5.0 100.0
Gemma2-2B-it 2.19 2.0 5.0 100.0
Qwen2.5-1.5B 1.43 1.0 4.0 35.9
Qwen2.5-1.5B-Instruct 1.43 1.0 4.0 35.9
Llama-3.1-8B 2.43 2.0 5.0 100.0
Llama-3.1-8B-Instruct 2.43 2.0 5.0 100.0
Table 6: Tokenization statistics across different models. Open-vocabulary models (GPT-2, Pythia-1.4B, Qwen2) require on average only 1.4–1.5 tokens per word, splitting roughly 36–42% of words into multiple tokens. By contrast, models with closed vocabularies (Gemma-2B, BERT variants, DeBERTa-v3-large) exhibit much finer subword granularity—averaging 2.2–3.1 tokens per word and splitting every word into multiple tokens. Median tokens per word reflect this: 1 for open-vocabulary models, 2–3 for closed-vocabulary models. Maximum tokens per word range from 4 (Qwen2) up to 8 (BERT).
Appendix C Linear Separability Analysis
Lexeme Inflection
Linear MLP Linear MLP
Model Slope R2 Slope R2 Slope R2 Slope R2
BERT-Base $-0.368$ $0.994$ $-0.376$ $0.924$ $0.011$ $0.063$ $0.013$ $0.096$
BERT-Large $-0.414$ $0.977$ $-0.331$ $0.878$ $0.007$ $0.025$ $0.005$ $0.020$
DeBERTa-v3-Large $-0.615$ $0.992$ $-0.527$ $0.935$ $-0.005$ $0.018$ $-0.003$ $0.005$
GPT-2-Small $-0.261$ $0.948$ $-0.049$ $0.006$ $-0.016$ $0.076$ $-0.012$ $0.032$
GPT-2-Large $-0.303$ $0.986$ $-0.389$ $0.860$ $-0.028$ $0.208$ $-0.017$ $0.183$
GPT-2-XL $-0.282$ $0.983$ $-0.367$ $0.868$ $-0.036$ $0.329$ $-0.022$ $0.315$
Pythia-6.9B $-0.213$ $0.992$ $-0.127$ $0.056$ $-0.009$ $0.053$ $-0.007$ $0.028$
Pythia-6.9B-Tulu $-0.195$ $0.988$ $-0.111$ $0.043$ $-0.005$ $0.016$ $-0.002$ $0.003$
OLMo-2-7B $-0.153$ $0.566$ $-0.186$ $0.614$ $0.011$ $0.067$ $0.015$ $0.137$
OLMo-2-7B-Instruct $-0.157$ $0.574$ $-0.178$ $0.615$ $0.009$ $0.051$ $0.013$ $0.106$
Gemma-2-2B $-0.161$ $0.413$ $-0.198$ $0.714$ $0.003$ $0.009$ $0.004$ $0.013$
Gemma-2-2B-Instruct $-0.116$ $0.233$ $-0.207$ $0.505$ $0.016$ $0.176$ $0.014$ $0.138$
Qwen2.5-1.5B $-0.260$ $0.629$ $0.040$ $0.007$ $0.004$ $0.007$ $0.015$ $0.054$
Qwen2.5-1.5B-Instruct $-0.191$ $0.392$ $0.126$ $0.077$ $0.002$ $0.002$ $0.015$ $0.050$
Llama-3.1-8B $-0.109$ $0.188$ $-0.163$ $0.387$ $0.015$ $0.105$ $0.014$ $0.101$
Llama-3.1-8B-Instruct $-0.076$ $0.103$ $-0.121$ $0.258$ $0.018$ $0.153$ $0.016$ $0.138$
Table 7: Regression slopes and R2 values for regressing probe accuracy on layers. Negative slopes indicate decreasing accuracy with layer depth.
@@ Appendix C Linear Separability Analysis
Table 8: Breakdown of inflection classification accuracy by morphological feature for each model, using linear regression classifiers. Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
@@
Table 9: Breakdown of inflection classification accuracy by morphological feature for each model, using MLP classifiers. Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
@@
Llama-3-8B 0.776 0.882 0.771 0.887 0.831 0.286 0.396 0.321 0.911
Llama-3-8B-Instruct 0.796 0.892 0.788 0.896 0.908 0.300 0.443 0.357 0.917
Table 10: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using linear regression classifiers. Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
@@
Llama-3-8B 0.798 0.879 0.807 0.886 0.800 0.214 0.679 0.393 0.882
Llama-3-8B-Instruct 0.824 0.893 0.826 0.895 0.831 0.257 0.689 0.429 0.887

### v4

@@ F.1 English Dataset Details
(a) Distribution of word categories
@@
(b) Distribution of inflection categories
Value
Average Words 6.5
Median Words 5
Minimum Words 1
Maximum Words 40
(c) Sentence length statistics
Table 4: Distribution statistics for the English dataset
F.2 Tokenization Statistics
@@
An important consideration for our analysis is how different models tokenize the words in our dataset. Table˜6 shows tokenization statistics across the models we analyze. Encoder-only models like BERT and DeBERTa tend to split words into more tokens than decoder-only models like GPT-2 and Qwen2, which may affect how information is encoded across layers.
Model Avg. tokens Med. tokens Max tokens Percent multitoken
per word per word per word
BERT variants 1.11 1.0 6.0 6.95
DeBERTa-v3-large 1.03 1.0 4.0 2.2
GPT-2 variants 1.52 1.0 5.0 42.25
Pythia-6.9B variants 1.48 1.0 5.0 39.1
OLMo2-7B variants 1.43 1.0 4.0 35.9
Gemma2-2B variants 1.19 1.0 4.0 16.55
Qwen2.5-1.5B variants 1.43 1.0 4.0 35.9
Llama-3.1-8B variants 1.43 1.0 4.0 35.85
Table 6: Tokenization statistics across different models (English only). Most models have an average of 1.0-1.5 tokens per word and a median of 1, indicating that most words are tokenized as a single unit. However, there is variation in the proportion of words split into multiple tokens. Decoder-only models (e.g., , GPT-2, Pythia, Qwen2, LLaMA) split 35-42% of words, while BERT and DeBERTa variants split fewer words (2-7%). Maximum tokens per word range from 4 to 6 across all models.
F.3 Effects of Tokenization
Figure 13: Effect of tokenization strategy on analogy completion rank. Each point corresponds to a model (color) and analogy (shape). The x-axis is the rank using whole-word representations. The y-axis is the rank using tokenized representations. Here, rank means the position of the expected word when all vocabulary words are sorted by similarity to the resulting embedding from vector arithmetic; lower is better. Points above the gray y=xy{=}x line mean tokenization hurts performance.
Tokenization is an essential component of language modeling. To test how tokenization influences our findings, we use analogy completion tasks in English (e.g., man:king::woman:?) and compare two approaches: averaging subtoken embeddings after standard tokenization and summing embeddings from whole-word tokens.
Whole-word representations markedly outperform averaged subtokens across all models (Figure˜13), implying that linguistic regularities are primarily stored in whole-word embeddings rather than compositionally across subtokens. Despite tokenization effects, our classifier results show consistent patterns across models using different tokenizers (see Table˜5), indicating robust encoding of lexical and morphological information.
Model HuggingFace ID
BERT-Base bert-base-uncased
BERT-Large bert-large-uncased
DeBERTa-v3-Large microsoft/deberta-v3-large
mT5-base google/mt5-base
GPT-2-Small openai-community/gpt2
GPT-2-Large openai-community/gpt2-large
GPT-2-XL openai-community/gpt2-xl
Pythia-6.9B EleutherAI/pythia-6.9b
Pythia-6.9B-Tulu allenai/open-instruct-pythia-6.9b-tulu
OLMo-2-1124-7B allenai/OLMo-2-1124-7B
OLMo-2-1124-7B-Instruct allenai/OLMo-2-1124-7B-Instruct
Gemma-2-2B google/gemma-2-2b
Gemma-2-2B-Instruct google/gemma-2-2b-it
Qwen2.5-1.5B Qwen/Qwen2.5-1.5B
Qwen2.5-1.5B-Instruct Qwen/Qwen2.5-1.5B-Instruct
Qwen2.5-7B Qwen/Qwen2.5-7B
Qwen2.5-7B-Instruct Qwen/Qwen2.5-7B-Instruct
Llama-3.1-8B meta-llama/Llama-3.1-8B
Llama-3.1-8B-Instruct meta-llama/Llama-3.1-8B-Instruct
Goldfish English goldfish-models/goldfish_eng_latn_1000mb
Goldfish Chinese goldfish-models/goldfish_zho_hans_1000mb
Goldfish German goldfish-models/goldfish_deu_latn_1000mb
Goldfish French goldfish-models/goldfish_fra_latn_1000mb
Goldfish Russian goldfish-models/goldfish_rus_cyrl_1000mb
Goldfish Turkish goldfish-models/goldfish_tur_latn_1000mb
Table 7: Canonical HuggingFace model IDs used to load models in our study.
Appendix G Attention Head Analysis
We conducted additional experiments analyzing attention head outputs alongside residual stream representations to understand how different components of transformer models contribute to linguistic encoding.
G.1 Methodology
We averaged activations across all attention heads at each layer for Qwen2.5-1.5B and Qwen2.5-1.5B-Instruct models using the English dataset. We then trained linear regression, MLP, and random forest classifiers on both attention head outputs and residual stream representations to compare their encoding patterns.
G.2 Results
Figure 14: Linguistic task accuracy for attention head outputs (solid lines) versus residual stream representations (dashed lines) across BERT and GPT-2 model families. Top row shows lemma accuracy, bottom row shows inflection accuracy. Columns represent different classifier types. Attention outputs show pronounced drops in lexical accuracy during middle layers, while residual streams maintain higher performance. Both streams maintain high inflectional accuracy, though residual streams consistently outperform attention outputs.
Figure 15: Classifier selectivity for attention head outputs (solid lines) versus residual stream representations (dashed lines) across BERT and GPT-2 model families. Top row shows lemma selectivity, bottom row shows inflection selectivity. Attention outputs show near-zero lemma selectivity throughout all layers, while inflection selectivity reaches moderate levels (0.4-0.5) in both streams, with residual streams showing slightly higher values.
Figure 16: Linguistic task accuracy for attention head outputs (solid lines) versus residual stream representations (dashed lines) across contemporary model families. Top row shows lemma accuracy, bottom row shows inflection accuracy. The pattern of attention outputs showing lower lexical accuracy in middle layers is consistent across all model families, while inflectional accuracy remains high in both streams.
Figure 17: Classifier selectivity for attention head outputs (solid lines) versus residual stream representations (dashed lines) across contemporary model families. The selectivity patterns mirror those seen in BERT and GPT-2 families, with attention outputs maintaining near-zero lemma selectivity and moderate inflection selectivity across all models.
The attention analysis reveals distinct encoding patterns between attention heads and residual streams. For lexical information, residual streams maintain relatively high accuracy (0.6-0.9) with gradual decline, while attention outputs show sharp drops to 0.2-0.4 in middle layers before modest recovery. For inflectional morphology, both components maintain high accuracy (0.7-1.0), though residual streams consistently outperform attention outputs, particularly in middle layers.
Selectivity analysis confirms that attention heads encode inflectional information more than lexical information, with inflection selectivity reaching 0.4-0.5 while lemma selectivity remains near zero. This supports the hypothesis that attention mechanisms primarily handle contextual relationships rather than parametric lexical knowledge.
Appendix H Steering Vector Analysis
We conducted steering vector experiments to test whether inflectional representations can be functionally manipulated and to understand model sensitivity to activation interventions.
H.1 Methodology
For each inflectional category, we computed steering vectors as:
$\mathbf{s}_{i}=\mathbf{\mu}_{i}-\lambda\cdot\frac{1}{|C|-1}\sum_{j\in C,j\neq i}\mathbf{\mu}_{j}$ (8)
We tested multiple values of $\lambda$ (1, 5, 10, 20, 100) and measured the impact on MLP classifier performance when adding these steering vectors to existing activations for 1000 test words. We evaluated both mean probability change and prediction flip rate across all models.
H.2 Results
Figure 18: Mean probability change for inflection prediction when applying steering vectors across different $\lambda$ values. Five panels show results for $\lambda\in\{1,5,10,20,100\}$. All models start with high effectiveness ($\approx$0.9-1.0) at layer 0. Most models maintain stable performance, but Qwen2.5 variants show pronounced sensitivity dips around 10% layer depth before recovering. Higher $\lambda$ values increase steering effectiveness while preserving the overall pattern.
Figure 19: Prediction flip rate when applying steering vectors across different $\lambda$ values. The flip rate patterns mirror the probability change results, with most models maintaining high rates (0.98-1.00) throughout all layers. Qwen2.5 variants show characteristic V-shaped dips to $\approx$0.60-0.70 around 10% layer depth. The consistency across $\lambda$ values suggests that steering effectiveness depends more on model architecture than intervention strength.
Steering vector analysis demonstrates that inflectional representations can be effectively manipulated across most models and layers. The majority of models show robust steering effectiveness (>0.95 mean probability change and flip rate) throughout all layers, indicating that inflectional information is functionally accessible and modifiable.
However, we observe model-specific sensitivity patterns. The Qwen2.5 variants show pronounced vulnerability in middle layers (around 10% depth), where steering effectiveness drops to 0.6-0.7 before recovering in later layers. This sensitivity pattern aligns with our intrinsic dimensionality findings, where these same models show extreme compression in middle layers (requiring only 1 component for 50-99% variance).
The consistency of steering effectiveness across different $\lambda$ values suggests that the underlying representational structure, rather than intervention magnitude, determines steering success. This supports our conclusion that models with compressed middle-layer representations are more susceptible to activation interventions.
Appendix I Random Forest Analysis
For our in-depth analysis of lexical information and inflectional morphology, we evaluated Random Forest classifiers as an additional non-linear baseline to complement our linear regression and MLP analyses.
I.1 Random Forest Classifier
$\hat{y}=\operatorname{argmax}_{j\in\{1,2,...,c\}}s_{j}(x)$ (9)
where $s_{j}(x)$ is the probability that the trained binary classifier assigns to class $j$ for the input xx [^5] We use scikit-learn’s implementation (Pedregosa et al., 2011) with the predict_proba() method to obtain probability scores.
Figure 20: Random Forest classifier performance across model layers for English. The left column shows accuracy for lemma (top) and inflection (bottom) prediction. The right column shows classifier selectivity (difference between linguistic and control task accuracy) for the same tasks. Each line represents a different model.
Figure 21: Random Forest classifier performance across languages and model layers. The left column shows accuracy for inflection prediction (lemma prediction omitted due to computational constraints with thousands of classes). The right column shows classifier selectivity.
I.2 English Results
Lexical Identity
Random Forest classifiers for lemma prediction were not comprehensively evaluated due to computational constraints arising from thousands of unique lemma classes. The one-vs-all training approach required for multi-class classification with Random Forests becomes prohibitively expensive with such a large number of classes.
Inflectional Morphology
Random Forest classifiers for inflectional morphology show markedly different patterns compared to linear regression and MLP classifiers (Figure˜20, bottom left). While linear and MLP classifiers maintain consistently high accuracy (0.9-1.0) across all layers, Random Forests exhibit lower overall accuracy (0.65-0.90) and a declining trend across layers. This contrasts sharply with the flat, high-performance curves observed for linear methods.
Selectivity Analysis
Random Forest selectivity for inflectional morphology (Figure˜20, bottom right) shows moderate values (0.2-0.4) but with more variability across models compared to linear methods. This suggests that while Random Forests can generalize beyond memorization, they are less consistent in their ability to extract the underlying morphological patterns compared to simpler linear approaches.
I.3 Cross-Linguistic Patterns
The multilingual Random Forest analysis (Figure˜21) reveals consistent cross-linguistic patterns while highlighting language-specific challenges.
Cross-Linguistic Performance
Random Forest classifiers for inflectional morphology show the same language ordering observed with linear and MLP classifiers: Russian and German achieve the highest accuracy (0.80-0.90), English and French show intermediate performance (0.70-0.85), Chinese maintains moderate accuracy (0.65-0.80), and Turkish consistently underperforms (0.55-0.75). This consistency across classifier types confirms that the observed patterns reflect genuine differences in how linguistic information is encoded rather than classifier-specific artifacts.
Selectivity Patterns
Random Forest selectivity patterns are similar to those of linear classifiers but with greater variance. Russian and German maintain the highest selectivity (0.3-0.5), while Turkish shows the lowest (0.1-0.3). This consistency suggests that the underlying representational structure, rather than classifier choice, drives the observed cross-linguistic differences.
Appendix J Classifier Error Analysis
We conducted a detailed error analysis of our classifiers to better understand their performance across different morphological features and languages. See Table˜8 through Table˜26 for the full results.
@@ I.3 Cross-Linguistic Patterns
Table 8: Breakdown of inflection classification accuracy by morphological feature for each model using linear regression classifiers (English). Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale. Comparative and superlative forms consistently show the lowest accuracy across all models, reflecting the challenges of these less frequent morphological categories.
@@
Table 9: Breakdown of inflection classification accuracy by morphological feature for each model using MLP classifiers (English). Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale. MLP classifiers provide modest improvements over linear regression, particularly for comparative and superlative forms, though the relative ordering across morphological features remains consistent.
@@
Llama-3.1-8B 0.776 0.882 0.771 0.887 0.831 0.286 0.396 0.321 0.911
Llama-3.1-8B-Instruct 0.796 0.892 0.788 0.896 0.908 0.300 0.443 0.357 0.917
Table 10: Breakdown of lemma classification accuracy by Part of Speech (POS) for each model using linear regression classifiers (English). Lemmas are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale. Performance varies significantly with frequency: frequent categories like nouns and verbs achieve higher accuracy, while infrequent categories like pronouns and prepositions show lower performance due to limited training examples.
@@
Llama-3.1-8B 0.798 0.879 0.807 0.886 0.800 0.214 0.679 0.393 0.882
Llama-3.1-8B-Instruct 0.824 0.893 0.826 0.895 0.831 0.257 0.689 0.429 0.887